2 men each working 5 hour per day can finish a job in 12 days. How long will 4 men each working 2 hours per day take to finish the same job?

Grade:12th pass

2 Answers

Pawan Prajapati
2 years ago
Our expert is working on this Class VIII Maths answer. We will update the answer very soon.
Jigyanshu Dixit
one year ago
let the number of days be x. The ratio of number of days=12:x Number of days it takes is inversely proportional to the number of hours each man works per day and the number of men at work. ∴ 12 : x = 4 × 2 : 2 × 5 ∴12:x=4×2:2×5 ⇒ 12 x = 8 10 ⇒ x = 10 × 12 8 = 15 ⇒12x=810⇒x=10×128=15 days ∴ ∴ 4 men each working 2 hours per day can finish the job in 15 days